OpenStudy (firejay5):

Simplify each Expression. 56. 3rd power of the square root of 5 * the square root of 5^3 58. 6th power of the square root of 81A^4B^8

OpenStudy (saifoo.khan):

\[\Large \sqrt[3]{5} \times \sqrt{5^3}\]this>

OpenStudy (firejay5):

Yes

OpenStudy (firejay5):

I need help on both please; I am confused and don't know what to do???

OpenStudy (saifoo.khan):

You can write that as: \[\Large 5^\frac13 \times 5^\frac32\] Taking base as common.. \[\Large 5^{\frac13 + \frac32}\]

OpenStudy (firejay5):

5 11/6

OpenStudy (firejay5):

is that right 5 ^\[5^ \frac{ 11 }{ 6 }\]

OpenStudy (firejay5):

@saifoo.khan

OpenStudy (saifoo.khan):

Yep.

OpenStudy (firejay5):

is it simplified

OpenStudy (firejay5):

I think it has to be in a radical

OpenStudy (saifoo.khan):

Now write it in radical form: 

OpenStudy (firejay5):

what about 58, I really don't get that one???

OpenStudy (saifoo.khan):

Try it out.

OpenStudy (firejay5):

well I would if I had some steps to do it

OpenStudy (firejay5):

could you help me on half of it

OpenStudy (firejay5):

it would mean a lot

OpenStudy (saifoo.khan):

Okay. start with factoring 81.

OpenStudy (firejay5):

81 has a square root of 9

OpenStudy (saifoo.khan):

Since the radical have the power 6. We should look for factor which can have group of SIX factors. Like 2*2*2*2*2*2. So we would've taken out 2. But now 81 has no such factors, so it will stay inside the radical. Do you agree till here?

OpenStudy (firejay5):

cause no number will multiply together to get 81

OpenStudy (saifoo.khan):

Exactly. So 81 will remain inside the radical. Next A. It have power of 4. Which is smaller than 6. So it will stay outside the radical too. Agree?

OpenStudy (firejay5):

it also had to be in radical answer as well

OpenStudy (firejay5):

A^4B^1 is what I got so far and the square root sign 81b^2

OpenStudy (firejay5):

\[A^4B^1\sqrt[6]{81B^2}\]

OpenStudy (saifoo.khan):

NO! A will remain inside the radical. Since it has power lesser than 6.\[\Huge \sqrt[6]{81A^4B^8}\]

OpenStudy (firejay5):

What about B you can take a pair of 6 with 2 b's leftover

OpenStudy (saifoo.khan):

YES! You're right. I was coming to B. \[\Large B \sqrt[6]{81A^4B^2}\]

OpenStudy (firejay5):

is that the simplified answer

OpenStudy (saifoo.khan):

yes sir.

OpenStudy (firejay5):

@saifoo.khan Here's what wolframalpha got 3^(2/3) (A^4 b^8)^(1/6)

OpenStudy (firejay5):

\[3^\frac{ 2 }{ 3 }\sqrt[6]{A^4B^8}\]